OpenStudy (darkigloo):

Given f(x)=sinx, use the tangent line at x=1 to estimate f(1.1). A. 0 B. 0.540 B. 0.841 D. 0.896 E. 1

OpenStudy (jdoe0001):

use the tangent line <--- meaning the derivative of the function

OpenStudy (darkigloo):

f '(x) = cosx

OpenStudy (jdoe0001):

yeap

OpenStudy (darkigloo):

so what do i do with that?

OpenStudy (jdoe0001):

so I gather.... you could just use f(1.1) = sin(1.1) and get it but you're asked to get it the same value, by means of the derivative thus f'(1) = cos(1)

OpenStudy (jdoe0001):

hmmm wait.. is not the same value

OpenStudy (anonymous):

The linear approximation of \(f\) at some point \(x_0\) is: \[ L(x) -f(x_0) = f'(x_0)(x-x_0) \]

OpenStudy (anonymous):

In this case, \(x_0 = 1\) and \(x = 1.1\).

OpenStudy (anonymous):

Do you get it, or do I have to continue?

OpenStudy (darkigloo):

so is this how it should look? L(1.1) - sin(1) = cos(1) (1.1-1)

OpenStudy (anonymous):

Yes, and they want you to solve for \(L(1.1)\approx f(1.1)\).

OpenStudy (anonymous):

Though to be honest, this question is a bit poor because there is nothing that makes \(\sin(1)\) any easier to calculate than \(\sin(1.1)\). It would have been better if they had you use \(\sin(\pi)\) to approximate \(\sin(3)\) or something.

OpenStudy (darkigloo):

oh ok. i got 0.896 as my answer.

OpenStudy (darkigloo):

ahh i understand. thanks